During this time of pandemic, many lending institutions offer a loan package to wage earners. Suppose loan company ABC offers you to borrow money amounting to PHP 50, 000 and payable in 4 years at a simple interest rate of 3% per annum, while loan company XYZ offers a 3% compounded annually. What loan company will you borrow money, then justify why you decide to choose to make loan with that particular loan company. thank uu so much for the help...
Answer by Boreal(15235)

first is 50000(1+(0.03/1))^4=56,275.44 PHP
second interest= 50000*0.03*4 the basic I=prt or $6000 interest, so the payback would be $56,000
The first is more expensive to do because it was compounded, so you were paying interest on interest as well.